OpenStudy (anonymous):

Factor completely: 8x2 + 26x + 6

OpenStudy (anonymous):

16+26x+6 2(8+13x+3)

OpenStudy (unklerhaukus):

\[8x^2 + 26x + 6\] \[8x^2 + 26x + 6=0\] \[x_{1,2}=\frac{-(26)\pm\sqrt{26^2-4(8)(6)}}{2(8)}\] \[(x-x_1)(x-x_2)=0\] \[(x-x_1)(x-x_2)\]

OpenStudy (anonymous):

not a choice

OpenStudy (anonymous):

\((4x+1)(2x+6)\)